OpenStudy (anonymous):

y-2 = -1/3(x + 6) write in standard form using integers :)

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Thank you c:

OpenStudy (anonymous):

is it multiple choice?

OpenStudy (anonymous):

so u can match my answer to the correct answer

OpenStudy (anonymous):

no :/ i have to create a math textbook and i have to work out the problems

OpenStudy (anonymous):

and idk how to solve this one

OpenStudy (anonymous):

\[\frac{ x }{ 3 }+y=0\]

OpenStudy (anonymous):

one moment

OpenStudy (anonymous):

ok no problem

OpenStudy (anonymous):

\[y-2=\frac{ -1 }{ 3}(x+6)\]

OpenStudy (anonymous):

actually the whole fraction is negative

OpenStudy (anonymous):

\[y-2=\frac{ -1 }{ 3}x-2\]

OpenStudy (anonymous):

it doesn't matter if i put the negative infront of the fraction, with the numerator, or with the denominator it still means that the fraction is negative.

OpenStudy (anonymous):

\[y=\frac{ -1 }{ 3}x\]

OpenStudy (anonymous):

added 1/3x to both sides to get x/3+y-0 which is the equation in standard form

OpenStudy (anonymous):

im so confused as to how you went from (x + 6) to -2

OpenStudy (anonymous):

\[\frac{ x }{3 }+y=0\]

OpenStudy (anonymous):

because i had to multiply -1/3 by x to get -1/3x and then i multiplied it by 6. -1/3 times 6 = -2.

OpenStudy (anonymous):

should i draw it?

OpenStudy (anonymous):

that would be great if you could

OpenStudy (anonymous):